OpenStudy (anonymous):

Multiply and simplify: radical 6 times the quantity radical 3+5 radical 2. A.3 square 3 +15square 2 B.3 square 2+5 square 12 C.3 square 2+10 square 3

OpenStudy (anonymous):

\[3\sqrt{3}+15\sqrt{2}\] \[3\sqrt{2}+5\sqrt{12}\] \[3\sqrt{2}+10\sqrt{3}\]

jhonyy9 (jhonyy9):

B is right